OpenStudy (anonymous):

Find the sum of 6c − 5c2 − 8 and 9c5 − c (A) 14c7 + 5c − 8 (B) 15c6 + 6c3 − 8 (C) 9c5- 5c2 + 5c − 8 (D) 9c5- 5c2 − 2

OpenStudy (anonymous):

(6c-5c2-8)+(9c5-c) =6c-5c2-8+9c5-c =9c5-5c2+6c-c-8 =9c5-5c2+5c-8